# Backwards Data-Flow Analysis using Prophecy Variable in the BuildIt System

### v1

@@ 5.2. Domain Specific Language Implementations
Many DSLs provide domain abstractions directly in the programming language to enhance developer productivity. Examples include visualization/graphics languages such as Tikz  (tikz), Vega/Vega-Lite (vega; vegalite), gnuplot (gnuplot), matplotlib (matplotlib), and Graphviz (graphviz); build system and build system generation languages such as make (make), cmake (cmake), and Ninja (ninja); and lexing and parsing languages like flex/lex bison/yacc (flex; yacc) and ANTLR (antlr). Because the main focus of these languages is increased developer productivity and not performance, implementations typically perform little to no forwards or backwards analysis to enable optimizations. Our research differs in that it uses prophecy variables for transformations and optimizations that require information about the future execution of the program.
@@
DSLs such as GraphIt (graphit; g2; ugf), TACO (taco), Halide (halide), Diderot (diderot), Taichi (taichi), SQL (mysql; sqlite) and others target domains in which performance is a primary concern. Implementations of these DSLs therefore often contain a range of both forward and backward program analyses to enable the generation of optimized code. Implementations typically feature standard programming language implementation components such as parsers, intermediate representations, and forwards and backwards analyses that operate on these intermediate representations. Our research differs in that it uses the lightweight BuildIt approach, which eliminates many of these components. The combination of prophecy variables and reexecution to deliver correct prophecy variable values enables BuildIt to implement transformations and optimizations that require information about the future execution of the program using only forward execution, with no backwards analysis and no intermediate representation.
@@
Partial evalution/multistage programming is an established language implementation technique with a long history of research (Futamura1971; JonesGomardSestoft1993; ConselDanvy1993; DaviesPfenning1996; TahaSheard1997; TahaSheard2000; CaretteKiselyovShan2009; RompfOdersky2010; Danvy1996; Taha2004). Examples of general purpose languages with multistage support include Template Haskell (sheard2002template), Lightweight Modular Staging in Scala (lms), Scheme (scheme), MetaML (TahaSheard2000), Meta OCaml, Mint (mint), Terra (terra; terra2) implemented in Lua, C macros, and C++ templates (vandevoorde2002c++). Multistage programming is often used to embed a DSL within a general purpose language. Examples include PyTorch (pytorch), Tensorflow (tensorflow), TVM (Chen2018TVM), AnyDSL (anydsl), Delite (delite), LMS (lms), BuilDSL (buildsl), JAX (jax2018github), SHiM (shim), and BREeze (tamara:meng-thesis:2023). Such language implementations often include sophisticated program analyses, either via standard programming language implementation techniques, or using history variables or their equivalent (the best of our knowledge, none of these systems use prophecy variables or their equivalent). Our research differs in that it 1) uses a lightweight approach with no intermediate representation and no program analysis of this (nonexistent) intermediate representation, 2) uses prophecy variables to enable transformations and optimizations that require information about the future execution of the program, and 3) uses repeated first stage executions to compute correct prophecy variable predictions, making it possible to implement transformations and optimizations that require information about the future execution of the program in a staged compilation system built on a standard programming language implementation that provides only forward program execution.
@@ 6. Conclusion
Supplemental Material for Backwards Data-Flow Analysis using Prophecy Variable in the BuildIt System

### v2

@@ 5.2. Domain Specific Language Implementations
Many DSLs provide domain abstractions directly in the programming language to enhance developer productivity. Examples include visualization/graphics languages such as Tikz  (Tantau, 2013), Vega/Vega-Lite (Contributors, 2025c, b), gnuplot (Williams and Kelley, 2011), matplotlib (Hunter, 2007), and Graphviz (Research, 2025); build system and build system generation languages such as make (Feldman, 1979), cmake (Inc., 2025), and Ninja (Martin, 2012); and lexing and parsing languages like flex/lex bison/yacc (Paxson, 2025; Johnson, 1975) and ANTLR (Parr, 2025). Because the main focus of these languages is increased developer productivity and not performance, implementations typically perform little to no forwards or backwards analysis to enable optimizations. Our research differs in that it uses prophecy variables for transformations and optimizations that require information about the future execution of the program.
@@
DSLs such as GraphIt (Zhang et al., 2018; Brahmakshatriya et al., 2021b, a), TACO (Kjolstad et al., 2018), Halide (Ragan-Kelley et al., 2013), Diderot (Chiw et al., 2012), Taichi (Hu et al., 2019), SQL (Widenius et al., 2002; Hipp, 2020) and others target domains in which performance is a primary concern. Implementations of these DSLs therefore often contain a range of both forward and backward program analyses to enable the generation of optimized code. Implementations typically feature standard programming language implementation components such as parsers, intermediate representations, and forwards and backwards analyses that operate on these intermediate representations. Our research differs in that it uses the lightweight BuildIt approach, which eliminates many of these components. The combination of prophecy variables and reexecution to deliver correct prophecy variable values enables BuildIt to implement transformations and optimizations that require information about the future execution of the program using only forward execution, with no backwards analysis and no intermediate representation.
@@
Partial evalution/multistage programming is an established language implementation technique with a long history of research (Futamura, 1971; Jones et al., 1993; Consel and Danvy, 1993; Davies and Pfenning, 1996; Taha and Sheard, 1997, 2000; Carette et al., 2009; Rompf and Odersky, 2010b; Danvy, 1996; Taha, 2004). Examples of general purpose languages with multistage support include Template Haskell (Sheard and Jones, 2002), Lightweight Modular Staging in Scala (Rompf and Odersky, 2010a), Scheme (Steele and Sussman, 1975), MetaML (Taha and Sheard, 2000), Meta OCaml, Mint (Westbrook et al., 2010), Terra (DeVito et al., 2013a, b) implemented in Lua, C macros, and C++ templates (Vandevoorde and Josuttis, 2002). Multistage programming is often used to embed a DSL within a general purpose language. Examples include PyTorch (Paszke et al., 2019), Tensorflow (TensorFlow Development Team, 2015), TVM (Chen et al., 2018), AnyDSL (Contributors, 2025a), Delite (Chafi et al., 2011), LMS (Rompf and Odersky, 2010a), BuilDSL (Brahmakshatriya and Amarasinghe, 2022), JAX (Bradbury et al., 2018), SHiM (Ray, 2023), and BREeze (Mitrovska, 2023). Such language implementations often include sophisticated program analyses, either via standard programming language implementation techniques, or using history variables or their equivalent (the best of our knowledge, none of these systems use prophecy variables or their equivalent). Our research differs in that it 1) uses a lightweight approach with no intermediate representation and no program analysis of this (nonexistent) intermediate representation, 2) uses prophecy variables to enable transformations and optimizations that require information about the future execution of the program, and 3) uses repeated first stage executions to compute correct prophecy variable predictions, making it possible to implement transformations and optimizations that require information about the future execution of the program in a staged compilation system built on a standard programming language implementation that provides only forward program execution.
@@ 6. Conclusion
Supplemental Material for Backwards Data-Flow Analysis using Prophecy Variables in the BuildIt System